Presented below is information related to Waterway Manufacturing Corporation.

Asset

Cost

Estimated Salvage

Estimated Life (in years)

A

$53,800 $6,100 10

B

32,200 5,200 9

C

36,500 4,100 9

D

19,000 1,500 7

E

24,000 3,300 6

(A) Compute the rate of depreciation per year to be applied to the plant assets under the composite method. (Round answer to 2 decimal place, e.g. 4.83%.)
(B) Prepare the adjusting entry necessary at the end of the year to record depreciation for the year.

(C)Prepare the entry to record the sale of asset D for cash of $5,400. It was used for 6 years, and depreciation was entered under the composite method.
Answer #1

SOLUTION

A. Depreciation = (Cost - Salvage value) / Estimated useful life

Depreciation of Asset A = (53,800 - 6,100) / 10 = 4,770

Depreciation of Asset B = (32,200 - 5,200) / 9 = 3,000

Depreciation of Asset C = (36,500 - 4,100) / 9 = 3,600

Depreciation of Asset D = (19,000 - 1,500) / 7 = 2,500.

Depreciation of Asset E = (24,000 - 3,300) / 6 = 3,450

Annual total depreciation = 4,770 + 3,000 + 3,600 + 2,500 + 3,450 = 17,320

Annual total cost = 53,800 + 32,200 + 36,500 + 19,000 + 24,000 = 165,500

Composite depreciation rate = Annual total depreciation / Annual total cost

= 17,320 / 165,500 = 10.47%

B. Journal entry

Accounts titles and Explanation Debit ($) Credit ($)
Depreciation expense 17,328
Accumulated depreciation 17,328

Depreciation = 165,500 * 10.47% =17,328

C.

Accounts titles and Explanation Debit ($) Credit ($)
Cash 5,400
Accumulated depreciation 13,600
Asset D 19,000